OpenStudy (anonymous):

help please! use integer values of x from –3 to 3 to graph the equation y = –2x2 + 3

OpenStudy (amistre64):

what is the value of –2x2 + 3, when x = -3 what is the value of –2x2 + 3, when x = -2 what is the value of –2x2 + 3, when x = -1 what is the value of –2x2 + 3, when x = 0 what is the value of –2x2 + 3, when x = 1 what is the value of –2x2 + 3, when x = 2 what is the value of –2x2 + 3, when x = 3 might want to keep track of those in a table of x and y values